# Visualizing LLM Latent Space Geometry Through Dimensionality Reduction

### v1

@@ 4.2 Large Norm of 0-th Sequence Position Latent States
Figure 4: Norm of latent states from intermediate layers from both GPT-2 and LLaMa along sequence positions. Norms were averaged over both samples and layers. The dataset used is PG-19.
@@
When performing dimensionality reduction using methods such as PCA or UMAP on a set of vectors (such as latent states), a few vectors with outlier norms can obscure the structures present between all other vectors in the set by overshadowing them [^1] In the UMAP case, this is applicable when the chosen metric is sensitive to norm outliers, such as with the Euclidean distance metric. Other metrics, such as cosine distance, would not be affected by norm outliers.. We analyze the intermediate layer norms of PG-19 hidden states averaged over the sample and layer dimensions for each sequence position. Figure 4 shows large spikes in the norm of the 0-th sequence position token for both LLaMa and GPT-2. This phenomenon has been noticed and analyzed in prior works and has functions which include acting as a bias term (sun2024massiveactivationslargelanguage).
As was similarly described by (sun2024massiveactivationslargelanguage), we find that these large spikes in latent norm occur in the LLaMa model at the initial token of a sequence. We find this interesting, as although the LLaMa model uses a <BOS> token at the start of all inputs (touvron2023llamaopenefficientfoundation; kudo-richardson-2018-sentencepiece), the spike in latent state norm for the 0-th sequence position is not limited to the appearance of the <BOS> token. Figure 5 shows histograms of the mean norm across intermediate layers for each vocab token in both GPT-2 and LLaMa when given as input to the model as the initial token. As can be seen, the vast majority of LLaMa vocab tokens take on large latent state norms when they are the initial token. We find this property intriguing as, unlike the learned positional embeddings of GPT-2, the relative RoPE positional encodings used by LLaMa do not provide the model with as trivial a way to determine whether an embedding belongs to a token in the 0-th sequence position. Nonetheless, tokens in the 0-th sequence position typically have a corresponding latent state in LLaMa with very high norm, indicating that LLaMa has likely developed a more complex mechanism to detect tokens at the 0-th sequence position.
@@
As was also described by (sun2024massiveactivationslargelanguage), the high-norm latent states emerge after a few initial Transformer layers and then reduce in norm in the final layers. Figure 6 shows the mean norm across all individual vocab tokens (separately input into the model) for each layer of both GPT-2 and LLaMa. As can be seen, the layers in the middle have very high norms, while those at the beginning and end do not.
Due to this observation of high latent state norms in the 0-th sequence position, in subsequent results, we often omit the latent states of the 0-th sequence position when performing dimensionality reduction and visualization so that it does not obscure all details in other latent states.
@@ 4.3 Layerwise Visualizations
Figure 7 shows the average norm of each layer of both GPT-2 and LLaMa across sequences of the PG-19 dataset. Initial token hidden states are excluded. In contrast with the consistently large norms seen in the initial token latent states of Figure 6, the norms of Figure 7 gradually increase across the layers and are never as large. Furthermore, there are patterns which are consistent across both GPT-2 and LLaMa, such as the large spike in norm in the last block of the Transformer and the drop in norm after the final norm layer.
@@
Figure 10: UMAP visualizations of GPT-2 and LLaMa pre-add intermediate layer latent states on the PG-19 dataset. For consistency with Figure 9, initial token hidden states are excluded.
@@ 4.5.1 GPT-2 Positional Embeddings
As has already been established in previous works, the positional embeddings of GPT-2 form a sort of “helix” in high-dimensional space (scherlis2022exploration; yedidia2023gpt2). This can clearly be seen in Figure 11. We also notice that a clear pattern across the sequence positions remains prominent for all combinations of PCA dimensions, demonstrating the high dimensionality of the geometric pattern formed by the GPT-2 positional embeddings.
@@ 4.5.2 LLaMa Positional Encodings
4.6 Repeating Token Experiment
We hypothesize that, in the case of RoPE-based models, the latent states of the same token repeated many times should eventually “converge” in latent space; after all, the attention heads of the model should view $2000$ previous tokens as little different from how it views $2001$ previous tokens. We experiment with inputs to LLaMa which are simply a sequence of the same token repeated many times. We only run this experiment for LLaMa as the same logic does not necessarily hold in the case of the learned positional embeddings of GPT-2, which introduces an absolute signature for each sequence position, whereas RoPE only affects relative distinctions between sequence positions.
We choose to repeat the token corresponding to the character “e” for the following reasons:
[…]
  • “e” is not commonly repeated in long sequences. At the same time, it is occasionally repeated in text, such as in "weeeeeeeeeeeeeeeee". Thus, “e” serves as a good “neutral” token in the sense that a sequence of repeating “e” tokens is neither very common nor extremely perplexing.
  • “e” by itself does not carry as much semantic information as other tokens which are complete words or concepts might. Of course, “e” can represent a mathematical constant; however, usually this would only become apparent when “e” is paired with other tokens of mathematical context.
We note that we tokenize the sequence of repeating “e” tokens with the <BOS> token as the initial token.
Figure 13: Latent state norm for each sequence position. The latent state norms were averaged across the intermediate layers.
Figure 13 shows the average latent state norm across the intermediate layers for each sequence position. After the initial spike of the initial <BOS> token, the norm of the repeating token latent states stays relatively low for $\sim 100$ tokens from gradually climbing to a large norm similar to that of the initial token latent states.
Figure 14: PCA of the repeating token latent states showing a convergence pattern. The latent states used were from the intermediate layers, and the latent states were averaged across the layers before dimensionality reduction or visualization.
Figure 14 shows a visualization after PCA of the repeating token latent states. As expected, a clear convergence pattern can be seen from the earlier to the later tokens. Notably, the initial (0-th) token is much closer to the final tokens than to the 1-st token. This is even true in the unit vector case, where only direction matters, and indicates that the later repeating “e” token latent states with high norm point in a similar direction to the initial <BOS> token. This is consistent with the findings of (sun2024massiveactivationslargelanguage), which found that “massive activations” (what we refer to as the latent states with huge norms) tend to be constant (all point in a similar direction).
@@ 5 Conclusion
We present a study into visualizing the internal representations of Transformer-based language models. Through systematic analysis of GPT-2 and LLaMa, we demonstrate how dimensionality reduction techniques can reveal geometric patterns that show how these models organize and process information. Our experiments highlight several notable phenomena. Most significantly, we identify a persistent geometric separation between attention and MLP component outputs (section 4.4), a pattern that appears consistent across different model architectures and has not been previously documented to our knowledge. We also noted the high norm of latent states at the initial sequence position (section 4.2), a phenomenon that extends beyond special tokens like <BOS> to many vocabulary tokens in LLaMa despite its use of relative position encodings. Additionally, we visualized the layerwise evolution of latent states (section 4.3), the geometric effects of sequence position (section 4.5), and experimented with repeated token sequences (section 4.6). Ultimately, we add to the growing body of interpretability research by motivating further work into analyzing feature geometry. Future work can help further our understanding of the dynamics of latent states within Transformer models across layers, models, and experimental conditions. By deepening our understanding of how these geometric structures emerge and behave, we move closer to principled, reliable interpretability methods capable of guiding the development of more transparent and understandable models.
@@ Acknowledgments
Intermediate Layers/Blocks — The middle portion of a Transformer, as defined in section 4.1. Section 4.1
Layer Dimension — The dimension of our generated dataset indexing Transformer’s latent captures along the model’s depth. Section 3.1
@@
Latent Space — The high-dimensional vector space in which latent states reside. Section 2.1
Learned Positional Encoding — A method of generating positional embeddings by learning weights. Section 2.4
@@
Norm — The L2 norm (Euclidean length/magnitude) of a latent vector. Section 4.2
@@
Positional Embedding — A positionally aware representation of token embeddings that encode absolute token order within the sequence. Section 2.4
@@
RoPE (Rotary Positional Encoding) — A method for generating positional embeddings by injecting position information through rotating attention query and key vectors in the attention mechanism according to token index. Section 2.4
@@
Sample Dimension — The dimension of our generated latent state dataset indexing latent states generated from different input samples. Section 3.1
Sequence Dimension — The dimension of our generated latent state dataset indexing token positions within a sequence. Section 3.1

### v2

@@ 4.2 Large Norm of 0-th Sequence Position Latent States
Figure 4: Norm of latent states from intermediate layers of both GPT-2 and LLaMa along sequence positions. Norms were averaged over both samples and layers. The dataset used is PG-19.
@@
When performing dimensionality reduction using methods such as PCA or UMAP on a set of vectors (such as latent states), a few vectors with outlier norms can obscure the structures present between all other vectors in the set by overshadowing them [^1] In the UMAP case, this is applicable when the chosen metric is sensitive to norm outliers, such as with the Euclidean distance metric. Other metrics, such as cosine distance, would not be affected by norm outliers.. We analyze the average intermediate layer norms of PG-19 latent states. Figure 4 shows large spikes in the norm of the 0-th sequence position token for both LLaMa and GPT-2. This phenomenon has been noticed and analyzed in prior works by Sun et al. in 2024 (sun2024massiveactivationslargelanguage) and has functions which include acting as a bias term.
As was similarly described by Sun et al. (sun2024massiveactivationslargelanguage), we find that these large spikes in latent norm occur in the LLaMa model at the initial token of a sequence. We find this interesting, as although the LLaMa model uses a <BOS> token at the start of all inputs (touvron2023llamaopenefficientfoundation; kudo-richardson-2018-sentencepiece), the spike in latent state norm for the 0-th sequence position is not limited to the appearance of the <BOS> token. Figure 5 shows histograms of the mean norm across intermediate layers for each vocab token in both GPT-2 and LLaMa when given as input to the model as the initial token. As can be seen, the vast majority of LLaMa vocab tokens take on large latent state norms when they are the initial token. We find this property intriguing as, unlike the learned positional embeddings of GPT-2, the relative RoPE positional encodings used by LLaMa do not provide the model with as trivial a way to determine whether an embedding belongs to a token in the 0-th sequence position. Nonetheless, tokens in the 0-th sequence position typically have a corresponding latent state in LLaMa with very high norm, indicating that LLaMa has likely developed a more complicated mechanism to detect tokens at the 0-th sequence position.
@@
As was also described by Sun et al. (sun2024massiveactivationslargelanguage), the high-norm latent states emerge after a few initial Transformer layers and then reduce in norm in the final layers. Figure 6 shows the mean norm across all individual vocab tokens (separately input into the model) for each layer of both GPT-2 and LLaMa. As can be seen, the layers in the middle have very high norms, while those at the beginning and end do not.
Due to this observation of high latent state norms in the 0-th (“initial”) sequence position, in subsequent results, we often omit the latent states of the 0-th sequence position when performing dimensionality reduction and visualization so that it does not obscure all details in other latent states.
@@ 4.3 Layerwise Visualizations
Figure 7 shows the average norm of each layer of both GPT-2 and LLaMa across sequences of the PG-19 dataset. Initial token latent states are excluded. In contrast with the consistently large norms seen in the initial token latent states of Figure 6, the norms of Figure 7 gradually increase across the layers and are never as large. Furthermore, there are patterns which are consistent across both GPT-2 and LLaMa, such as the large spike in norm in the last block of the Transformer and the drop in norm after the final norm layer.
@@
Figure 10: UMAP visualizations of GPT-2 and LLaMa pre-add intermediate layer latent states on the PG-19 dataset. For consistency with Figure 9, initial token latent states are excluded.
@@ 4.5.1 GPT-2 Positional Embeddings
As has already been established in previous works by Scherlis in 2022 and Yedidia in 2023 (scherlis2022exploration; yedidia2023gpt2), the positional embeddings of GPT-2 form a sort of “helix” in high-dimensional space. This can clearly be seen in Figure 11. We also notice that a clear pattern across the sequence positions remains prominent for all combinations of PCA dimensions, demonstrating the high dimensionality of the geometric pattern formed by the GPT-2 positional embeddings.
@@ 5 Conclusion
We present a study on visualizing the internal representations of Transformer-based language models. Through systematic analysis of GPT-2 and LLaMa, we demonstrate how dimensionality reduction techniques can reveal geometric patterns that show how these models organize and process information. Our experiments highlight several notable phenomena. Most significantly, we identify a persistent geometric separation between attention and MLP component outputs (section 4.4), a pattern that appears consistent across the tested model architectures and has not been previously documented to our knowledge. We also noted the high norm of latent states at the initial sequence position (section 4.2), a phenomenon that extends beyond special tokens like <BOS> to many vocabulary tokens in LLaMa despite its use of relative position encodings. Additionally, we visualized the layerwise evolution of latent states (section 4.3) and the geometric effects of sequence position (section 4.5). Ultimately, we add to the growing body of interpretability research by motivating further work into analyzing feature geometry. Future work can help further our understanding of the dynamics of latent states within Transformer models across layers, models, and experimental conditions. By deepening our understanding of how these geometric structures emerge and behave, we move closer to principled, reliable interpretability methods capable of guiding the development of more transparent and understandable models.
@@ Acknowledgments
Intermediate Layers/Blocks — The middle portion of a Transformer, as defined in the Experimental Setup section. Section 4.1
Layer Dimension — The dimension of our generated dataset that indexes different layers/components across the Transformer’s depth. Section 3.1
@@
Latent Space — The high-dimensional vector space in which internal model representations (latent states) reside. Section 2.1
Learned Positional Embeddings — A method of generating positional embeddings by learning high-dimensional additive embedding vectors. Section 2.4
@@
Norm — The L2 norm (Euclidean length/magnitude) of a vector. Section 4.2
@@
RoPE (Rotary Positional Encoding) — A method for injecting position information through rotating attention query and key vectors in the attention mechanism. Section 2.4
@@
Sample Dimension — The dimension of our generated latent state dataset that indexes different input samples. Section 3.1
Sequence Dimension — The dimension of our generated latent state dataset that indexes different token positions within a sequence. Section 3.1